If d2x/dt2 + g(x-a)/b = 0, (a,b,g being positive constants) and x=a' ans dx/dt = 0 when t = 0, show that x = a + (a'-a) cos[(√g)t/b].


1
Expert's answer
2021-06-02T17:19:29-0400

given

"\frac{d^2x}{st^2}+\frac{g}{b}(x-a)=0; \space x=a' ,\frac{dx}{dt}=0 \space when \space t=0
\\"

we can write it

"\frac{d^2x}{dt^2}+\frac{g}{b}x-\frac{ag}{b}=0\space \\
\frac{d^2x}{dt^2}+\frac{g}{b}x=\frac{ag}{b}\space \\" .........[1]

Complementary Function

 Auxiliary equation is:

"m^2+\frac{g}{b}=0
\\m^2=-\frac{g}{b}\\
m=\pm\sqrt{\frac{g}{b}}"

thus the C.F. solution is

"x = e^{0x}(Acos(\sqrt{\frac{g}{b}}t + Bsin(\sqrt{\frac{g}{b}}t))\\
x =(Acos(\sqrt{\frac{g}{b}}t + Bsin(\sqrt{\frac{g}{b}}t))\\" ........................[CF]

Particular Solution

"\frac{d^2x}{dt^2}+\frac{g}{b}x=f(t)\space \\
where \space f(t)=\frac{ag}{b}\space \\"

So, we should probably look for a solution of the form:

x=C........[2]

Where the constant C is to be determined by direct substitution and comparison:

Differentiating [2] wrt x twice we get:

x′ =0

x''=0

Substituting these results into the DE [1] we get:

"(0) + \frac{gc}{b} = \frac{ag}{b}\\ C=a"

thus the P.I. solution is

x=a ....................................[PI]


General Solution = CF + PI

"x =(Acos(\sqrt{\frac{g}{b}}t + Bsin(\sqrt{\frac{g}{b}}t))+a\\"

Next we apply the initial conditions:

"x=a' ,\frac{dx}{dt}=0 \space when \space t=0"

So that:

a'=A cos(0)+B sin(0)+a

A=a'−a

And differentiating the above result wrt t

"x'(t)=−A\sqrt{\frac{g}{b}}sin(\sqrt{\frac{g}{b}}t)+B\sqrt{\frac{g}{b}}cos(\sqrt{\frac{g}{b}}t)"

And again applying the initial conditions:

"0=−A\sqrt{\frac{g}{b}}sin0+B\sqrt{\frac{g}{b}}cos0\\
B=0"

And so the complete solution is:


"x
(
t
)
=
(
a
'
−
a
)
cos
(
\sqrt{\frac{g}{b}}
t
)
+
a"